OpenStudy (cherrilyn):

with reduction formula and techniques, evaluate the integral of cos^2(sint)costdt

OpenStudy (anonymous):

let u = cost and du = sint ^_^ that's all, give it a try now

OpenStudy (cherrilyn):

thanks! I'll try it now

OpenStudy (cherrilyn):

If u = cost and du=sint...what should I do with the cos^2?

OpenStudy (anonymous):

u = sint, du =cost dt. Then from here it becomes cos^2(u) du. Which turns into cos(u)*cos(u) which you can integrate with integration by parts.

OpenStudy (anonymous):

good luck :) np

OpenStudy (anonymous):

cos^2 = u^2 :)

OpenStudy (anonymous):

^_^ substitute cos x with u

OpenStudy (cherrilyn):

so u = cos x not cos t?

OpenStudy (anonymous):

lol cost :)

OpenStudy (anonymous):

here : \[\int\limits \cos^2 t (cost) (sint) dt =-\int\limits u^3 du \rightarrow = - \frac{u^4}{4} + c \] \[= \frac{\cos^4 t}{4} + c\] better ? :)

OpenStudy (anonymous):

I forgot the minus in the last step, add it ^_^

OpenStudy (anonymous):

remember u = cos t, so du= - sin t

OpenStudy (cherrilyn):

yes! thank you so much :)

OpenStudy (anonymous):

np ^_^